investing $17985 on march 1st 2021, i pay .02% per month and the invester wants their money on January , 2024. how much will I give the invester back


1
Expert's answer
2021-03-04T10:33:15-0500

"solution\\

principal= 17,985\ dollars\\
payment\ rate\ per\ month=.02\ percent\\
payment \ duration=12+12+10=34 \ months\\"


"\$17,985\times {(1+0.0002)}^t\\
\$17.985\times ({1.0002})^t\\
t=34\\
\$17,985\times 1.0068=\$18,107.298\\"